OpenStudy (artfis):

(1+w) (4w - 5) = 0 My last question tonight.... :)

OpenStudy (jtvatsim):

So, you have two sets of parentheses that multiply to give you 0. If you think about it the only way you can multiply two numbers together and get 0 is if at least one of them is 0. So, we set each parentheses equal to 0 and solve both for w: 1 + w = 0 OR 4w - 3 = 0

OpenStudy (artfis):

Then that would leave..... W = -1 or W = 4/3

OpenStudy (jtvatsim):

sorry you are right, but I used the wrong number in mine -.- you had 4w - 5 = 0, so w should be 5/4.

OpenStudy (artfis):

Haha it's all good. I see it now. Thanks so much. Medal for you standing up. CHYEAH!

OpenStudy (jtvatsim):

good night! :)